OpenStudy (anonymous):

A circle is defined by the equation x2 + y2 – x – 2y – = 0. What are the coordinates for the center of the circle and the length of the radius?

OpenStudy (xapproachesinfinity):

use this formula (x-a)^2 + (y-b)^2=r^2 you will need to complete the square to ge the answer

OpenStudy (anonymous):

oh there is suposed to be a fraction after the 2y that didnt show up. after the 2y it is -11/4 as a fraction

OpenStudy (xapproachesinfinity):

it doesn't matter just use the formula, the fraction put it on the right hand side

OpenStudy (xapproachesinfinity):

when completing the square there will be more number to add or substract from both side to

OpenStudy (anonymous):

what do i put in for the variables.

OpenStudy (xapproachesinfinity):

first group x^2 with -x, y^2 with -2y x^2-x+y^2-2y=11/4 x^-x+1/4+y^2-2y+1-1/4-1=11/4 (x-1/2)^2+(y-1)^2=11/4+1+1/4

OpenStudy (xapproachesinfinity):

continue

OpenStudy (xapproachesinfinity):

if we have x^2-bx we complete the square like this x^2-bx+(b/2)^2-(b/2)^2=(x-b/2)^2-(b/2)^2

OpenStudy (anonymous):

well as i was working it out i ended up with \[x ^{2}+y ^{2}=1.75\]

OpenStudy (xapproachesinfinity):

No I already gave you half answer the center should be (1/2, 1) the radius is 2

OpenStudy (anonymous):

oh i see where i messed up. thank you for the help

OpenStudy (xapproachesinfinity):

You're welcome